# Gravity-aware Grasp Generation with Implicit Grasp Mode Selection for Underactuated Hands

### v2

@@ V-B Simulation Benchmark
[…]
@@
Even though some existing works evaluated their performance against various geoemetric properties (such as shape), there are limited examples that evaluate against various physical properties such as object weight or friction coefficient. In this work, we propose to evaluate the grasp against a varying object weight, since in the light-weight region, it captures whether the hand can reach the object, while in the heavy-weight region, it captures whether the hand can lift the object. Note that the capability to lift heavier objects is equivalent to the capability to lift the same objects with less grasp force, which is critical for handling fragile objects.
Figure 8 plots the SR and CR against the object weight from 0.1 kg to 1.5 kg. In the light-weight region (less than 0.5 kg), Ours-Precision performed similarly with VGN while VN-EdgeGraspNet outperformed both. This is reasonable because we share the same backbone as VGN, and the author of VN-EdgeGraspNet reported that their grasp representation is more advantageous. In the heavy-weight region (more than 1 kg), however, Ours-Precision outperformed VGN and performed on par with VN-EdgeGraspNet. This indicates that even limited to precision grasping, our gravity-rejection score is beneficial in grasping heavy objects. One hypothesis is that if the hand only supports precision grasping, introducing the gravigy-rejection score to the edge-grasp representation will provide a better score. Nevertheless, our focus is on leveraging power grasp. Ours, with all features including power grasp annotation, significantly outperformed the others, with 20.9 % improvement for both SR and CR in the case of 0.1 kg and 41.9 / 48.7 % improvement of SR / CR in the case of 1.5 kg compared with VN-EdgeGraspNet. Its superiority in the light-weight region indicates that power grasping is robust against positioning error, and that in the heavy-weight region indicates that it’s robust against gravity.
@@ V-C Comparison on the Rotation Representation
Following Qin et al. [2], we also select rotation matrix as the rotation representation due to its continuity in the SO(3). However, [2] lacks a discussion on how to select the basis vectors of the rotation matrix, i.e., how we should define the TCP coordinate. A rotation matrix $R=[\bm{e}_{x},\bm{e}_{y},\bm{e}_{z}]\in\mathbb{R}^{3\times 3}$ is constructed by three basis vectors. The network estimates two of them $\tilde{\bm{e}}_{x},\tilde{\bm{e}}_{z}$ and reconstructs $R$ by:
@@ V-D Validation with a Physical System
In this subsection, we validate our approach with a physical system, as shown in Fig. 9. Since we only have physical access to the Robotiq hand, we trained a baseline version of our model to mimic [5] by (i) training data only contained precision grasps, (ii) we only used grasp validness and ignored gravity-rejection score for the inference, (iii) quaternion rotation representation. We selected nine kinds of objects from the YCB dataset [27], which we didn’t use for training, and created 20 scenes with 51 objects in total in a simulator. In the real system, we overlaid the scene’s mesh with the real-time-streamed point cloud and manually adjusted the objects’ pose in the real world. Depth image was acquired from a pair of RGB cameras and a learning-based stereo inference [28]. Table II summarizes the result, showing a significant improvement in our approach over the baseline. Note that we cannot compare the absolute SR / CR with the simulation evaluation in V-B due to the sim-to-real gap. Nevertheless, we can observe a consistent result that our approach outperforms the baselines in both environments.
Figure 9: Target objects (left) and capture (middle, right) of the validation with a physical system.
TABLE II: Success ratio (SR) and clear ratio (CR) in the real robot validation
@@
VI Conclusion
In this paper, we presented a data generation and learning framework to leverage power grasping in addition to precision grasping. We proposed to train a neural network against the continuous gravity-rejection score, that is the magnitude of disturbance in the gravity direction a grasp can support, rather than the grasp probability, in order to let the network prioritize power grasping while still keeping precision grasping as a secondary choice. We also proposed a data generation pipeline to efficiently create a dataset with gravity-rejection score annotation. We provided a qualitative visualization and analysis of the network output and showed the effect of the combination of gravity-rejection score head and grasp validness head. For the quantitative evaluation, we proposed to evaluate the performance against varying object weights. A simulation evaluation proved a significant improvement in our approach, where power grasping improves the grasp robustness in both positioning accuracy error and gravity direction force disturbance. Finally, we provided real robot validation, which showed that our approach is also effective in a physical system.

### v3

@@ V-B Simulation Benchmark
In this subsection, we compare our approach (ours) with two baselines: Breyer et al. [4] (VGN) since we share the same backbone, and Huang et al. [7] (VN-EdgeGraspNet) since they reported state-of-the-art performance. While both [4] and [7] provide a pre-trained weight for Franka Emika’s hand [30] 111[7] performed a real-robot validation with the Robotiq gripper in their paper, but they only made Franka Emika version’s weight public. , we target Robotiq 2F-85 [14] hand; thus, we created a simulation model for both hands with an identical grasp force and contact parameters [^2] We set 40 N for grasp force, 0.75/0.5 for static/dynamic friction coefficient, 5 for Hunt Crossley dissipation, 108 for hydroelastic-modulus.. As the simulator, we used Drake [29] since its hydroelastic contact model [28] allows for a stable simulation of power grasping by hands with a passive parallel-link mechanism. To clarify the effect of power grasps, we also trained our model (Ours-Precision) for Franka Emika’s hand with only precision grasp annotation in the training data.
@@
While SR and CR are intuitive metrics, they mix two distinct aspects of the grasp quality: positional accuracy and robustness against disturbance. We, therefore, propose to evaluate grasp detectors against varying object weights since, in the lightweight region, they mainly capture positional accuracy, while in the heavy region, they capture both aspects. The ability to support a heavier object also means a grasp can lift the same objects with less grasp force, which is critical for handling fragile objects. Note that changing other physical properties, such as the friction coefficient, is another option, but it is harder to disentangle the two aspects. For example, lower friction requires both high positional accuracy (because only highly antipodal grasps are acceptable in precision grasping) and high robustness (because the object will slip more easily).
Figure 8 plots the SR and CR against the object weight from 0.1 kg to 1.5 kg. In the light-weight region (less than 0.5 kg), Ours-Precision performed similarly with VGN while VN-EdgeGraspNet outperformed both. This is reasonable because we share the same backbone as VGN, and the author of VN-EdgeGraspNet reported that their grasp representation is more advantageous. In the heavy-weight region (more than 1 kg), however, Ours-Precision outperformed VGN and performed on par with VN-EdgeGraspNet. This indicates that even when limited to precision grasping, our gravity-rejection score is beneficial in grasping heavy objects. One hypothesis is that if the hand only supports precision grasping, introducing the gravigy-rejection score to the edge-grasp representation will provide a better score. Nevertheless, our focus is on leveraging power grasping. Ours, with all features including power grasp annotation, significantly outperformed the others, with 20.9 % improvement for both SR and CR in the case of 0.1 kg and 41.9 / 48.7 % improvement of SR / CR in the case of 1.5 kg compared with VN-EdgeGraspNet.
It was surprising that even in the light-weight region, Ours outperformed the other cases. One reason is simply because the Robotiq 2F-85 hand has a slightly larger fingertip than that of the Franka Emika’s hand, but we also observed that power grasping is more robust against position error. In a power grasp, the palm or the proximal finger link hits the object first, then the hand closes, and the distal fingers wrap around it. Even if the hand is inserted too deeply, the grasp is still valid as long as the object doesn’t fall down in the case of a side grasp. Conversely, even when the hand insertion is too shallow, the distal fingers’ wrap-around motion still “pulls” the object into the hand, or, in the worst case, results in a precision grasp.
Figure 8 also shows that Ours had the minimum drop of SR / CR due to the increasing object weight. This quantitatively proves the benefit of power grasping. Qualitatively, we observed that power grasping frequently occurred for cylinder-shaped objects and that it was especially effective in lifting heavy cylinders laid down on the surface by a top-down grasp. We also observed that when the hand attempts to perform a power grasp on a thin object laid down on the surface, the grasp typically fails due to the fingertip’s collision with the surface and the finger’s undesired displacement due to passive compliance. However, such failure mode was rare, and in most cases, a precision grasp was selected. This shows that utilizing the grasp validness is effective in allowing the network to automatically fall back to precision grasping when power grasping is not available.
@@ V-C Comparison on the Rotation Representation
Following Qin et al. [1], we also select the rotation matrix as the rotation representation due to its continuity in the SO(3) [31]. However, [1] lacks a discussion on how to select the basis vectors of the rotation matrix, i.e., how we should define the TCP coordinate. A rotation matrix $R=[\bm{e}_{x},\bm{e}_{y},\bm{e}_{z}]\in\mathbb{R}^{3\times 3}$ is constructed by three basis vectors. The network estimates two of them $\tilde{\bm{e}}_{x},\tilde{\bm{e}}_{z}$ and reconstructs $R$ by:
@@ V-D Validation with a Physical System
In this subsection, we validate our approach with a physical system, as shown in Fig. 9. Since we only have physical access to the Robotiq hand, we trained a baseline version of our model to mimic [4] by (i) training data containing only precision grasps, (ii) we only used grasp validness and ignored gravity-rejection score for the inference, (iii) quaternion rotation representation. We selected nine kinds of objects from the YCB dataset [32], which we didn’t use for training, and created 20 scenes with 51 objects in total in a simulator. In [4], the objects were put in a box and dumped on the table. However, since our target objects were often rigid and round-shaped, they frequently rolled out of the workspace. Thus, we first created the scenes with the simulator, overlaid the scene’s mesh with the real-time-streamed point cloud, and manually adjusted the objects’ poses in the real world. Depth image was acquired from a pair of RGB cameras and a learning-based stereo inference [33]. Table II summarizes the result.
The absolute number of SR / CR was lower than the value reported by [4]. We assume this is because of the different object properties. In the case of [4], most objects were light and soft toys, which are relatively easy to grasp. Indeed, our objects also contained such an object (061_foam_brick), and we observed that in any case, the grasp succeeds. However, our objects contained more heavy and rigid objects. For example, the unopened 005_tomato_soup_can weighed 350 g, and its hard metal exterior made the contact region small. The object even harder to grasp was 035_power_drill, with more than 600 g weight and unevenly distributed mass. Other factors that may have contributed to the difference include the camera system, hand-eye calibration, and the underlying robot controller. Similarly, due to the sim-to-real gap, we cannot compare the absolute SR / CR with the simulation evaluation in V-B.
In order to avoid such biases, our evaluation was performed under the exact same system configuration, and our approach significantly outperformed the baseline. This result is consistent with the simulation and further justifies the effect of our contribution.
Figure 9: Target objects (left) and capture (middle, right) of the validation.
TABLE II: Success ratio and clear ratio in the real robot validation
@@
VI Limitation
In this work, we reconstruct the hand pose through the location of the voxel corresponding to the TCP and the regressed rotation. This representation, though, is fragile against the rotation regression error, since a small error in the rotation results in a large error in the contact regions. We frequently observed failure cases in which the hand approach direction was inaccurate, causing the hand to grasp at an improper position (sometimes even not making contact with the object at all). Another major failure case was when the finger collided with the object during the approaching motion due to an incorrect hand yaw rotation. Our attempt in Sec. V-C is one countermeasure for such a problem, but there is still a large margin for improvement. Contact-based grasp representations [3, 7] are more robust to rotation error. Finally, it is difficult to generate power grasps due to the inconsistent contact region among different grasp modes, even though our gravity-rejection score is also applicable to them. An approach to detect power grasps through contact-based representations remains in our future work.
VII Conclusion
In this paper, we presented a data generation and learning framework to leverage power grasping in addition to precision grasping. We proposed to train a neural network to predict a continuous gravity-rejection score which is the magnitude of disturbance in the gravity direction that a grasp can support. Unlike the grasp probability, this allows the network prioritize power grasping while still keeping precision grasping as a secondary choice. We also proposed a data generation pipeline to efficiently create a dataset with gravity-rejection score annotation. We provided a qualitative analysis of the network output and showed the effect of gravity-rejection score in generating power grasps. For the quantitative evaluation, we proposed to evaluate the performance against varying object weights. A simulation evaluation proved a significant improvement in our approach, where power grasping improves the grasp robustness to positioning accuracy error and gravity direction force disturbance. Finally, we provided real robot validation, which showed that our approach is also effective in a physical system.